An amount of R5 500 is invested for a period of 4 years in a savings account. For the first year, the investment grows at a simple interest rate of 11% p.a. and then at a rate of 12,5% p.a. compounded quarterly for the rest of the period. Determine the value of the investment at the end of the 4 years.


1
Expert's answer
2021-07-09T14:15:57-0400

P = R5500

t = 4


For the first year, the investment grows at 11% simple interest

"\therefore FV= PV (1+rn)"

Where "r=11\%" and "n=1"

"FV=5500(1+(1\u00d70.11))"

"FV=R6105"

At the beginning of the next year till the end of the year, t=3, the amount is compounded quarterly.

Given as

"FV=PV (1+\frac{r}{m})^{nm}"


PV= R6105 ; r= 12.5%

n=3 ;. m=4

"FV=6105(1+\frac{0.125}{4})^{3×4}"

"FV=R8831.88"


The value of the investment after 4 years is R8831.88 .